OpenStudy (anonymous):

which of the following vectors has a direction of 311? a) <7, -7 square root3> b) <2, -5> c) <2 square root 3, -4> d) <-1, square root 3> e) none of these

OpenStudy (anonymous):

plzz help

OpenStudy (johnweldon1993):

Direction can be found using \[\large \tan^{-1}(\frac{y}{x})\]

OpenStudy (anonymous):

how can we find 311 then?

OpenStudy (johnweldon1993):

Well if we know that \[\large \theta = \tan^{-1}(\frac{y}{x})\] We know that \(\large \theta = 311\) so we just need to see which coordinates will make our equation = 311 So for example the first choice you have there...x = 7 and y = -7√3 so plug those in...and see if it comes out to 311

OpenStudy (anonymous):

thanks it helps a lot